OpenStudy (anonymous):

Given a = 2, b = -2, and k = 3, what is the equation of the graph if the parent function is y= square root x

OpenStudy (anonymous):

@jim_thompson5910 can you give me what to solve agaom?

jimthompson5910 (jim_thompson5910):

the general form is \[\Large y = a\sqrt{bx-h} + k\]

OpenStudy (anonymous):

2√(-2x)-h+3?

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

\[\Large y = a\sqrt{bx-h} + k\] \[\Large y = 2\sqrt{-2x-0} + 3\] \[\Large y = 2\sqrt{-2x} + 3\] so you were close